OpenStudy (anonymous):

A ship is 2.8 degrees off course. If the ship is traveling at 14.0 miles per hour, how far off course will it be after 2 hours?

OpenStudy (itsstribe):

If it's going 14mph, it will have traveled 28 miles in 2 hours. s = d/t, so d = st sin 2.8° = x/28 x = 28 sin 2.8° x approx = 28 * 0.04885 = 1.3678 It'll be approx 1.3678 miles off course.